# Ego-to-Exo: Interfacing Third Person Visuals from Egocentric Views in Real-time for Improved ROV Teleoperation

### v3

@@ IV-B Proof of Concept: 2D Indoor Navigation
Experimental setup. The proof-of-concept experiments are conducted with TurtleBot4, a 2D ground robot that can be teleoperated with egocentric views from its front-facing monocular camera. It has only two degrees of freedom (DOF) for linear and angular velocity - which simplifies the motion kinematics for tracking its instantaneous position and orientation. We teleoperate it to collect visual data with a USB camera at p resolution in office, laboratory, and hallway scenarios. The experiments are designed to validate the proposed algorithm by evaluating ground plane estimation and reprojection errors.
@@
Geometric validation: reprojection error analysis. We first evaluate the reprojection errors of known reference points from the generated Ego-to-Exo views and the estimated ROV pose. As shown Fig. 3, we use standard checkerboard corners as reference points from egocentric views and then evaluate the reprojection errors for those points from exocentric views. This test is iterated over different sets of past egocentric images, each corresponding to a different EOB distance. As shown in Fig. 3(a), a checkerboard is viewed from different EOB distances (further back into the past) indicated by the parameter . More specifically, is the number of frames between the current egocentric view and the selected EOB view. The corresponding reprojection error is plotted in Fig. 3(b), which shows how the estimation is accurate for lower values of , and gradually degenerates for . This is consistent with our visual observation of the projected ROV point cloud, i.e., it is on the ground plane with accurate orientation based on the SLAM trajectory estimates.
@@
Geometric validation: ground plane estimation. We also visualize the accuracy of the estimated ground plane in these experiments for qualitative assessments. Fig. 3(b) illustrates two cases: one for with a low reprojection error, and another for with a high error. As seen, the estimated ground plane (and the drawn sample cube) validates the geometric accuracy for the case. On the other hand, a misaligned ground plane for the case demonstrates the underlying error in pose estimation as well as in the reprojection process. Essentially, the geometric accuracy of our proposed Ego-to-Exo algorithm depends on the trajectory estimation performance of the SLAM system. More importantly, a teleoperator can make the right viewpoint selection in real-time for optimal Ego-to-exo rendering.
@@
Computational efficiency. We analyze the computational complexity of the Ego-to-Exo algorithm for different configurations to ensure real-time execution in resource-constrained edge devices onboard standard ROV platforms. Table I shows the memory requirement of our algorithm for different choices of buffer size. The memory footprint is less than MB for a buffer size of up to frames, making it highly efficient. Table II demonstrates that the end-to-end framework maintains a consistent output rate of over FPS (frames per second), making it suitable for integration in existing teleoperation engines. In fact, our algorithm only adds overhead on the SLAM backbone for the global map update.
| Buffer size (# of frames) | |||||
@@
| Method | SLAM Only | SLAM + Ego-to-Exo |
| Output image rate | FPS | FPS |
| Map update rate | FPS | FPS |
IV-C Field Deployment: 3D Underwater Cave Exploration
@@
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV performs full 6-DOF motions. While the motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as a guide [49]. The mission objective is to navigate the ROV - feet deep inside the cave through its complex structures, and then safely return it to the surface. The videos are recorded at p resolution with a GoPro11 camera mounted on BlueROV2 and then compressed to p the within the Ego-to-Exo framework. In addition to evaluating the geometric accuracy and robustness, we consider how informative the generated Ego-to-Exo views are compared to traditional consoles for ROV teleoperation.
@@
Real-time map update and teleoperation. In addition to the Ego-to-Exo view generation and ROV pose rendering, our framework simultaneously updates a 3D map with extracted feature points from the SLAM system. Fig. 4 shows an ROV’s trajectory mapped during an underwater cave mission in Devil’s Springs, Florida. As seen, the generated Ego-to-Exo views embed significantly more peripheral information about the scene. The exocentric view of the ROV pose and its relative distance from cave walls or overhead obstacles are useful to surface operators for obstacle avoidance and efficient decision-making. Additionally, the 3D map shows the ROV’s past trajectory and its current pose which are useful to analyze the mission progress, which is not possible in traditional teleoperation consoles. Such a global view of the trajectory is also useful during emergency evacuation and recovery. Beyond cave exploration, these features will be crucial in ROV-based subsea surveillance and search-and-rescue operations as well.
@@
Accuracy and robustness. Due to the complex scene geometry and absence of a smooth ground plane inside underwater caves, we adopt a homography estimation approach for the performance validation. As shown in Fig. 5, April-Tag [50] corners are used as reference points for reprojection. Specifically, we compute the homography transformation between the egocentric and synthesized exocentric views to visualize the reprojection errors. We use a sample 2D logo and project it onto the reference April-Tag surface using the homographic transform. The unskewed planar projection validates the accuracy of the Ego-to-Exo pose estimation and point cloud rendering processes.
| # | Question | mean, std |
| System Usability Scale (SUS)[51] | ||
@@
Observations: strengths and limitations. Our experiments reveal some key strengths of the proposed teleoperation framework. First, the generated exocentric views closely resemble the actual EOB views during a smooth trajectory, which is usually the case for subsea exploration and surveying tasks. Second, the buffer memory works as a backup during a temporary failure of the SLAM system, typically observed at turning corners or due to abrupt motion. In such cases, our algorithm retains historical poses from its buffer memory; teleoperators can utilize this for situational awareness to safely anchor or pause the mission until communication is restored. On the other hand, its heavy dependency on the SLAM backbone leads to some inherent limitations. Feature-based monocular SLAM systems often fail in feature-deprived noisy underwater scenes, which leads to inaccurate pose tracking and thus inaccurate Ego-to-Exo view synthesis. Tracking 6-DOF ROV motion from monocular vision is particularly challenging with no additional sensor to recover the scale information [52, 53]. We observe some instances where the estimated ROV pose is incorrectly scaled in the rendering. To address this, multi-sensor fusion-based underwater SLAM backbones [54, 55] can be utilized in more critical applications.
@@ IV-D Subjective User Study
The user study is conducted with multiple underwater cave exploration data. A BlueROV2 recorded egocentric video feeds inside the caves at up to -meter penetrations. Later on, human participants, between the ages of - with no/little teleoperation experience, evaluate the ease of operation with our developed console and compare it to traditional consoles. Their feedback is recorded using the System Usability Scale (SUS) [51], with our interface achieving an average SUS score of . We also formulate an independent set of questions that reflect the teleoperator’s preference for the novel features of our method. The individual questions and corresponding scores are presented in Table III. Some key observations from this study are listed below.
1.
@@
The operators report a significantly less workload (perceived cognitive load) in conducting complex tasks such as object following and structure mapping.
V Improved Underwater ROV Teleoperation
@@
Multiple augmented viewpoints. We validate the utility of our proposed Ego-to-Exo teleoperation interface through further experiments on underwater cave exploration data. Our expedition in cave segments at Devil’s Springs, Florida reveals that when ROVs move slowly against strong currents, extending the exocentric viewpoint distance can significantly improve teleoperation. This is achieved by tuning the queue parameters , , and in the proposed TeleOp interface. We consistently find that exocentric views are more informative, especially for about - seconds preceding the ROV position during navigation. The multiple preceding views offered by our interface are particularly useful for mapping large structures such as newly discovered cave segments or shipwrecks. As Fig. 6 shows, the synthesized viewpoints provide more spatial context, enabling operators to control the ROV efficiently around complex underwater structures.
@@
Efficient teleoperation in complex missions. We also conducted trials inside an underwater cave system at Orange Grove, Florida, and a grotto system at Hudson, Florida. We observe that maneuvering the robot by following the caveline with egocentric views is challenging because little/no ambient light penetrates inside underwater caves. Despite using powerful lights, problems such as moving shadows and scattered waves create significant blind spots. Consequently, tracking and following the caveline or any other navigation markers [49] without any peripheral view is extremely disorienting to the operator. In some cases, we observe that the cavelines get blended with the texture and features of cave walls in noisy conditions; see Fig. 7. In such scenarios, shifting the viewpoint to exocentric views allows easier identification of cavelines against the surrounding and overhead cave walls. Additionally, the augmented 3D map displays the robot’s pose, allowing much safer maneuvering of the vehicle to its desired orientation.
@@
Safer navigation in hazy low-light conditions. Underwater caves present a unique formation of silt and sediment on their floor that results from erosion over extended periods. The silt is susceptible to disturbance from external factors, such as the motion of underwater ROVs or the turbulence generated by their propellers. Although ROV operators pay close attention to avoid contact with floor and cave-walls, it is often unavoidable due to buoyancy imbalance and strong flow of water. Dislodging the sediments results in cloudy or hazy conditions that obscure visibility. Bright lights from the ROV reflect from these suspended particles and make it even more challenging to capture clear imagery of the surroundings. In such cases, third-person EOB views from behind the ROV offer a clearer and more informative perspective for navigation; see Fig. 7. It improves spatial awareness and helps the operator to safely move away from the sediment formations toward open, accessible areas and avoid obstructing other scuba divers in the process.
VI Conclusion & Future Work

### v5

@@ IV-B Proof of Concept: 2D Indoor Navigation
Experimental setup. The proof-of-concept experiments are conducted with TurtleBot4, a 2D ground robot that can be teleoperated with egocentric views from its front-facing monocular camera. It has only two degrees of freedom (DOF) for linear and angular velocity, which simplifies the motion kinematics for tracking its instantaneous position and orientation. We teleoperate it to collect visual data with a USB camera at p resolution in office, laboratory, and hallway scenarios. The experiments are designed to validate the proposed algorithm by evaluating ground plane estimation and reprojection errors.
@@
Geometric validation: reprojection error analysis. We first evaluate the reprojection errors of known reference points from the generated EgoExo views and the estimated ROV pose. We use standard checkerboard corners as reference points from egocentric views and then evaluate the reprojection errors for those points from exocentric views. This test is iterated over different sets of past egocentric images, each corresponding to a different EOB distance. As shown in Figure 3(a), a checkerboard is viewed from different EOB distances (further back into the past), indicated by the parameter . More specifically, is the number of frames between the current egocentric view and the selected EOB view. The corresponding reprojection error is plotted in Figure 3(b), which shows how the estimation is accurate for lower values of , and gradually degenerates for . This is consistent with our visual observation of the projected ROV point cloud, i.e., it is on the ground plane with accurate orientation based on the SLAM trajectory estimates.
| # Ego | Inlier | Plane | Normal | Altitude |
| Frames | Fraction () | RMSE () | Drift () | Drift () |
Geometric validation: ground plane estimation. We adopt four metrics to evaluate the quality of ground plane estimation: inlier fraction, plane RMSE error, temporal drift in plane normal, and temporal drift in altitude. The inlier fraction for each frame reports the ratio of inliers to total candidate points () after RANSAC fit:
where are the fitted plane parameters, are the candidate 3D points, and is the distance threshold (see Eqn. 3). A higher indicates that the majority of candidate points are consistent with a single ground plane. Subsequently, the point-to-plane distances of inliers are calculated to quantify the fitting residual as root-mean-square error (RMSE):
A lower RMSE reflects a tighter fit around the estimated plane. Next, to assess temporal consistency, we compute the angular difference between consecutive plane normals:
The mean angular drift across frames is reported, where low values indicate temporal stability. Finally, the altitude at instance is computed as the vertical distance of the camera center to the estimated plane:
In the absence of true measurement, the computed (scaled) altitude is not meaningful; however, a low deviation across frames indicates that the synthesized 2.5D ground remains consistent for visualization.
In 2D indoor setup, the robot’s camera is rigidly mounted with negligible roll and pitch variation, so the estimated ground-plane normal is expected to align with the camera’s vertical axis and remain stable across frames. Consequently, the plane inlier fraction should be consistently high, and the residual error should approach zero. The results obtained from several trials in office, laboratory, and hallway scenarios are summarized in Table I. A high inlier fraction and low normal drift across frames confirm the robustness of our approach under such structured conditions; please refer to the next section for further evaluation in unstructured settings.
@@
Figure 3(b) illustrates two representative examples from an office scene: one for with a low reprojection error, and another for with a high error. As seen, the estimated ground plane normal validates the geometric accuracy for the case. On the other hand, a misaligned plane normal for the case demonstrates the underlying error in pose estimation as well as in the reprojection process. Essentially, the geometric accuracy of our proposed algorithm depends on the pose estimation performance of the SLAM system.
@@
Computational efficiency. We analyze the computational complexity of the proposed algorithm for different configurations to ensure real-time execution in resource-constrained edge devices onboard standard ROV platforms. Table II shows the memory requirement of our algorithm for different choices of buffer size. The memory footprint is less than MB for a buffer size of up to frames, making it highly efficient. Table III demonstrates that the EgoExo framework maintains a consistent output rate of over FPS (frames per second). The added computation for ground estimation slightly reduces the scene update rate in EgoExo++, but still maintains over FPS, making it suitable for integration in existing teleoperation engines.
| Buffer (# of Frames) | |||||
@@
| Method | SLAM | SLAM+EgoExo | SLAM+EgoExo++ |
| Exo Image | FPS | FPS | FPS |
| Scene Update | FPS | FPS | FPS |
IV-C Field Deployment: 3D Underwater Caves
@@
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV performs full 6-DOF motions. While the motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as guides [9]. The mission objective is to navigate the ROV - feet deep inside the cave through its complex structures, and then safely return it to the surface. The videos are recorded at p resolution with a GoPro11 camera mounted on BlueROV2 and then compressed to p within the EgoExo framework. In addition to evaluating the geometric accuracy, we consider how informative the generated views are compared to traditional consoles for ROV teleoperation.
@@
Real-time map update and teleoperation. In addition to the exocentric view generation and ROV pose rendering, the EgoExo framework simultaneously updates a sparse map with extracted feature points from the SLAM system. Figure 4 shows an ROV’s trajectory mapped during our trial in an underwater cave in Peacock Springs, Florida. As seen, the generated EgoExo views embed significantly more peripheral information about the scene. The exocentric view of the ROV pose and its relative distance from cave walls or overhead obstacles are useful to surface operators for obstacle avoidance and efficient decision-making. Additionally, the 3D map shows the ROV’s past trajectory and its current pose, which are useful to analyze the mission progress, which is not possible in traditional teleoperation consoles. Such a global view of the trajectory is also useful during emergency evacuation and recovery. Beyond cave exploration, these features will be crucial in ROV-based subsea surveillance and search-and-rescue operations as well.
@@
Validation: homographic projection. Due to the complex scene geometry inside underwater caves, we adopt a homography estimation approach for the performance validation. As shown in Figure 5, April-Tag [66] corners are used as reference points for reprojection. Specifically, we compute the homography transformation between the egocentric and synthesized exocentric views to visualize the reprojection errors. We use a sample 2D logo and project it onto the reference April-Tag surface using the homographic transform. The unskewed planar projection validates the accuracy of the EgoExo pose estimation and point cloud rendering processes.
Validation: ground plane and 2.5D view. The ground plane estimation in the field is assessed following the same method as the indoor validation; Table IV summarizes the results for trials performed at two different cave systems. In addition to the four metrics defined earlier in Eqn. 8- 11, a success rate is also reported, since a valid ground plane cannot always be recovered in unstructured cave scenes. The success rate is defined as the percentage of frames in which a plane can be reliably fitted from the sparse SLAM feature points.
| Field Trials | Peacock Springs | Devil’s Springs |
| Ego Frames | segments | segments |
| Success Rate () | ||
| Inlier Fraction () | ||
| Plane RMSE () | ||
| Normal Drift () | ||
| Altitude Drift () | ||
The results in Table IV show that although trials in Devil’s Springs cave systems have a higher success rate in detecting the ground plane, the plane quality from Peacock Springs cave systems is consistently better. This difference can be attributed to the more complex cave structure and the challenging ROV trajectory executed in the latter case. In Devil’s Springs, several obstacles (e.g., large rocks) appeared directly in front of the ROV, forcing the operator to ascend and maneuver around. The terrain itself had high altitude variations, composed of rocks, boulders, and scattered pebbles, in contrast to the relatively smooth sedimentary floor observed in Peacock Springs. The reconstructed ground map from Peacock Springs shows consistent elevation and orientation (see Figure 4(b)), supporting the quantitative results. More snapshots from the two sites are provided in Figure 6. As seen, the rocky terrain in Devil’s Springs and the resulting jerky vehicle motion led to larger errors in ground plane estimation, greater deviations in the fitted normal, and higher variability in estimated altitude.
Observations: strengths and limitations. Our experiments reveal some key strengths of the proposed teleoperation framework. First, the generated exocentric views closely resemble the actual EOB views during a smooth trajectory, which is usually the case for subsea exploration and surveying tasks. Second, the buffer memory works as a backup during a temporary failure of the SLAM system, typically observed at turning corners or due to abrupt motion. In such cases, our algorithm retains historical poses from its buffer memory; teleoperators can utilize this for situational awareness to safely anchor or pause the mission until communication is restored. On the other hand, its heavy dependency on the SLAM backbone leads to some inherent limitations. Feature-based monocular SLAM systems often fail in feature-deprived, noisy underwater scenes, which leads to inaccurate pose tracking and thus inaccurate EgoExo view synthesis. Tracking 6-DOF ROV motion from monocular vision is particularly challenging with no additional sensor to recover the scale information [67, 68]. We observe some instances where the estimated ROV pose is incorrectly scaled in the rendering. To address this, multi-sensor fusion-based underwater SLAM backbones [69, 70] can be utilized in more critical applications.
| # | Questions | Mean, Std. Dev. |
@@ IV-D Subjective User Study
The user study is conducted with multiple underwater cave exploration data collected during our field trials. A BlueROV2 recorded egocentric video feeds inside the caves at up to -meter penetrations. Later on, human participants, between the ages of - with little/no prior teleoperation experiences, evaluate the ease of operation with our developed console and compare it to traditional consoles. Their feedback is recorded using the System Usability Scale (SUS) [71], with our interface achieving an average SUS score of . We also formulate an independent set of questions on the teleoperator’s preference for the novel features of our method. The individual questions and corresponding scores are presented in Table V. Some key observations from this study are listed below.
1)
@@
The operators report a significantly lower workload (perceived cognitive load) in conducting complex tasks such as object following and structure mapping.
V Improved Underwater ROV Teleoperation: Strengths, Challenges, and Limitations
@@
Multiple augmented viewpoints. We validate the utility of our proposed EgoExo teleoperation interface through further experiments on underwater cave exploration data. Our expedition in cave segments at Devil’s Springs, Florida, reveals that when ROVs move slowly against strong currents, extending the exocentric viewpoint distance can significantly improve teleoperation. This is achieved by tuning the queue parameters , , and in the proposed TeleOp interface. We consistently find that exocentric views are more informative, especially for about - seconds preceding the ROV position during navigation. The multiple preceding views offered by our interface are particularly useful for mapping large structures such as newly discovered cave segments or shipwrecks [72, 73]. As Figure 6 shows, the synthesized viewpoints provide more spatial context, enabling operators to control the ROV efficiently around complex underwater structures.
2D and 2.5D exocentric view. Our EgoExo pipeline synthesizes third-person views as flat 2D projections of the robot model onto a reference egocentric view from the past. The EgoExo++ advances from this purely image-based rendering to a 2.5D representation by recovering and texturing a dense ground surface in 3D space. This addition provides both altitude awareness and geometric context relative to the terrain, enabling operators to maintain safe clearance above uneven ground [38]. Insights from our prior user study also emphasized the value of adjustable viewpoints, such as bird’s-eye or side perspectives. While earlier EOB viewpoints offered multiple exocentric views, they remained anchored to the robot’s trajectory and fixed reference frames. The EgoExo++ view is no longer tied to a fixed reference image; the virtual viewpoint can be freely adjusted, offering better situational awareness during ROV teleoperation [25].
@@
Efficient teleoperation in complex missions. We conducted extensive field trials across multiple underwater cave systems, including Orange Grove, Devil’s Springs, and Peacock Springs, as well as inside a grotto system in Hudson, Florida. We observe that maneuvering the robot by following the caveline with egocentric views is challenging because little/no ambient light penetrates inside underwater caves. Despite using powerful lights, problems such as moving shadows and scattered waves create significant blind spots [74]. Consequently, tracking and following the caveline or any other navigation markers [9] without any peripheral view is extremely disorienting to the operator. In some cases, we observe that the cavelines get blended with the texture and features of cave walls in noisy conditions; see Figure 7. In such scenarios, shifting the viewpoint to exocentric views allows easier identification of cavelines against the surrounding and overhead cave walls. Additionally, the augmented 3D map displays the robot’s pose, allowing much safer maneuvering of the vehicle to its desired orientation [75].
@@
Safer navigation in hazy low-light conditions. Underwater caves present a unique formation of silt and sediment on their floor that results from erosion over extended periods. The silt is susceptible to disturbance from external factors [76], such as the motion of underwater ROVs or the turbulence generated by their propellers. Although ROV operators pay close attention to avoid contact with the floor and cave walls, it is often unavoidable due to buoyancy imbalance and strong flow of water. Dislodging the sediments results in cloudy or hazy conditions that obscure visibility [77]. Bright lights from the ROV reflect from these suspended particles and make it even more challenging to capture clear imagery of the surroundings. In such cases, third-person EOB views from behind the ROV offer a clearer and more informative perspective for navigation, as shown in Figure 7. It improves spatial awareness and helps the operator to safely move away from the sediment formations toward open, accessible areas and avoid obstructing other scuba divers in the process [78, 9].
Operator-ROV shared autonomy. In subsea teleoperation, collaborative decision-making frameworks split responsibilities so that humans set high-level goals while the vehicle plans and autonomously executes low-level actions [13]. EgoExo++ augmented visuals strengthen this “human-in-the-loop” paradigm by providing an interactive shared scene where operator intent (e.g., safe altitude, keep-out zones) can be directly expressed and the ROV autonomously confirms execution. For instance, the 2.5D view helps the operator determine a safe ground clearance, which the ROV can then autonomously maintain. Additionally, higher-level human-robot interactions can be integrated in the EgoExo++ interface: the operator can draw an intended path directly on the exocentric view [79], then the ROV can plan and follow an optimal path accordingly. Beyond navigation, augmented visuals have high demand in shared telemanipulation tasks such as delicate object grasping [80], valve control [81], artifact collection [46], etc. Our proposed views provide the operator with critical situational cues in such tasks. For instance, a side-view perspective helps position the ROV with respect to the target, then the operator can switch to a close-up egocentric view for precise manipulation [13, 82]. Overall, EgoExo++ serves as a shared perceptual layer: it enhances situational awareness with explicit geometric cues and allows the operator to specify intent more precisely, promising a safer and more effective teleoperation and telemanipulation.
Digital Twins and shadows. A digital shadow creates a virtual replica of the robot and its environment, enabling operators to practice missions, rehearse manipulation tasks, and refine control strategies [83]. While a shadow is a passive replica, a digital twin offers a bidirectional data pipeline and predictive simulation, thereby closing the feedback loop [84]. In this context, EgoExo++ complements DT-based training by providing geometrically consistent exocentric perspectives and interactive Real2Sim 2.5D reconstructions derived from mission SLAM data. During rehearsal, such views allow operators to anticipate spatial challenges, practice navigation in cluttered cave-like terrains, and visualize manipulators reaching the target from third-person perspectives [85]. EgoExo++ views can also be rendered on HMIs, where identical head motions may be mapped to different outcomes depending on the selected visualization mode [86]. For instance, a head tilt in egocentric mode can directly control the ROV body, whereas the same movement in exocentric mode can control the virtual camera viewpoint, with no impact on the ROV. Rehearsing such multi-visual feedback and control mappings in high-fidelity simulator engines will significantly improve operator skills in high-risk, time-critical missions.
Challenges in ground estimation. Our field trials in diverse underwater caves and grotto systems reveal that the irregular and deceptive terrain poses several unique challenges in estimating the ground surface. As illustrated in Figure 8, elevation slope or sharp jumps may fall outside the fitting capability of plane detection algorithms, leading to fragmented or distorted ground estimation. Additionally, narrow passages create occlusions and limited visibility of the ground, causing gaps in both feature tracking and surface mapping. The presence of large protruding structures that appear ground-like in texture can lead to incorrect segmentation of the actual ground surface. These issues collectively challenge our EgoExo++ pipeline, occasionally resulting in an unreliable representation of the terrain.
VI Conclusion and Future Work